OpenStudy (anonymous):

Find dy/dx if y=ln(x^2+5x) My teacher said it was (2x+5)/(x^2+5x) But I don't know how to get the answer. Please help me.

OpenStudy (campbell_st):

well when you differentiate a log function \[y = \ln(f(x)\] the result is a fraction, the derivative over the function \[\frac{dy}{dx} = \frac{f'(x)}{f(x)}\] so in your question the function is \[f(x) = x^2 + 5x\] then the derivative is \[f'(x) = 2x + 5\]